OpenStudy (anonymous):

Finish the calculation below by arranging the steps to solve for x. x= 24 6x=2x+20 4x=20 x=5 6x=2x+28 x=16 x=7 4x=28 step 1: 6x-4=2x+24 step 2: step 3: step 4: i will give medal

OpenStudy (anonymous):

Oh okay I think I get it now so step 2: 6x=2x+20 Step 3: 4x=20 Step 4: x= 5

OpenStudy (er.mohd.amir):

6x-4=2x+24 6x-2x=24+4 4x=28 x=28/4 x=7 when side change of variable or constant then its sign change + to - and - to +.

OpenStudy (anonymous):

I'm pretty sure there is more than one correct answer.

OpenStudy (er.mohd.amir):

if equation linear then solution only one and if quadratic then solution is 2. linear eq has power of varible that is x is one.